OpenStudy (anonymous):

ABCD is a square where A is the point (0,2) and C is the point (8,4). AC and BD are diagonals of the square and they intersect at E. Find the coordinates of B and D.

OpenStudy (anonymous):

I have found out the coordinate of E which is(4,3) and the equation of Bd and length of AE

OpenStudy (amistre64):

slope is 1/4?

OpenStudy (amistre64):

so perp slope is -4

OpenStudy (anonymous):

the ans for the coordinate of B is(5,-1) & D is (3,7)

OpenStudy (anonymous):

but i do not know how to derive at it

OpenStudy (amistre64):

the slope form a to c aint 1 so its not a square that situated nicely like that is it?

OpenStudy (amistre64):

slope from a to c is: 2/8 = 1/4

OpenStudy (anonymous):

ABCD is a square in the question

OpenStudy (amistre64):

e is a+c/2 for midpoint 8,4 0,2 ---- 8,6 /2 = 4,3 e is good

OpenStudy (amistre64):

y=-4x+4(4)+3 y=-4x+19 is the equation of the line perp to ac and thru 4,3

OpenStudy (amistre64):

hmmm, so. ; this laptop doesnt make this easier lol

OpenStudy (amistre64):

we need a vector that is the length of ae

OpenStudy (anonymous):

its \[\sqrt{17}\] length of ae

OpenStudy (amistre64):

8,4 -4,3 ---- <4,1>; mag = sqrt(17) that ougt to be fun to play with

OpenStudy (anonymous):

i cant find a way to find the coordinate through the length of ae LOL

OpenStudy (amistre64):

so the vector representation of say eb is our -4/1 slope

OpenStudy (amistre64):

<1,-4> is out vector ... duh lol

OpenStudy (amistre64):

form e move 1,-4 and -1,4 to get to the other corners

OpenStudy (anonymous):

i was tinking of finding the coordinate through the x-axis difference and y-axis difference but it doesnt seem to work

OpenStudy (amistre64):

4,3 1,-4 ----- 5,-1 4,3 -1,4 ----- 3,7

OpenStudy (anonymous):

i get it thx:) alot

OpenStudy (amistre64):

took a bit my my brain clicked lol

hero (hero):

So what were the correct points? For some reason, I'm getting (3,7) and (5,-1), but the question is which one is B and which one is D

OpenStudy (anonymous):

Hi im very sorry could you explain what is a vector?

OpenStudy (amistre64):



OpenStudy (amistre64):

a vector can be represented as a directed line segment; something like an arrow with distance and direction defines a vector

OpenStudy (anonymous):

how did u arrive at a vector of (-1,4)

OpenStudy (amistre64):



OpenStudy (amistre64):

the slope of a line IS its vector i took the slope of the line from A to C

OpenStudy (amistre64):

not drawn correctly in this pic, but same concept nonetheless

OpenStudy (anonymous):

gradient AC= 1/4 thus the vector is (-1,4 ) and gradient AB= 4 vector is (-4,1)?

OpenStudy (amistre64):

the slope of the line from A to C is: C (8,4) -A (0,2) -------- 8, 2 ; slope = 2/8; 1/4 the vector from A to C is then <4 , 1>

OpenStudy (amistre64):

a vector is notated the same as a point, except for the < > parts that indicate it as a vector. it is 4 wide and 1 tall; the components of its gradient (slope)

OpenStudy (anonymous):

How did u get C (8,4) -A (0,2) -------- 8, 2

OpenStudy (amistre64):

its called subtraction

OpenStudy (amistre64):

8-0 , 4-2 these are the parts of the slope formula

OpenStudy (amistre64):

its just easier for me to do them in this way then to try and sort out what goes where when its already done

OpenStudy (amistre64):

but to be clear about the vector from A to C, I was a little off in my explaining. The vector from A to C is actually <8,2>; half of this is going to be to the midpoint of E, since a midpoint is in the middle <8,2> ------ = <4,1>, is our vector from A to E 2

OpenStudy (anonymous):

i see thank you i will try to understand & do the question first, if I still do not understand could I still look for you? thx::)